iosangel:

A phone that was worth $499 last year is worth $299 this year. What is the percent of decrease? Assuming the percent of decrease remains the same, what equation models the value of the phone any number of years after this year?

iosangel:

I got the first part which is 40% but the second part im not sure

iosangel:

i got the formula but im stuck again y=b(1-r)^t and i plug but i don't know what is b y=b(1-0.4)^t

justjm:

b is just the initial value, so you started at 499.

iosangel:

the original price like this? y=499(1-0.4)^t

iosangel:

nvm i dont think thats right

justjm:

It's correct, because each year, the value is decreasing by a percent, not by a constant value. It's not linear, it's exponential.

iosangel:

I got it wrong tho

Nnesha:

" Assuming the percent of decrease remains the same, what equation models the value of the phone any number of years after `this year`?"

iosangel:

oh 299 lol

iosangel:

Thank you

Nnesha:

:) good job keep it up!

justjm:

As you can see, sometimes small words have big consequences